OpenStudy (anonymous):

A teacher set up groups of three students to work together. She asked them to solve the equation independently, then check their answers with each other. Here are the results: x divided by three minus two-fifths equals negative one-fifth Chet: x equals negative three-fifths Lauren: x equals one and four-fifths Bradley: x equals three-fifths Which student found with the correct answer? Chet Lauren Bradley None of them

OpenStudy (anonymous):

Can you put the words into numbers? its more better in numbers.

OpenStudy (anonymous):

x/3 -2/5 = -1/5 x/3=1/5 x=3/5 which would make Bradley correct

OpenStudy (anonymous):

x/3 - 2/5 = -1/5 --- add 2/5 to each side x/3 = -1/5 + 2/5 --- combine like terms x/3 = 1/5 --- multiply entire equation by 3 x = 1/5(3) x = 3/5

OpenStudy (anonymous):

I agree with @ilovemathbutneedhelp ...bradley is correct